OpenStudy (anonymous):

A can of white paint costs $16 while a can of green paint costs $20. Susan is mixing 12 cans of paint and will pay $19 per can for the mixture. How many cans of white paint will she purchase?

OpenStudy (anonymous):

A) 9 B) 6 C) 1 D) 3

OpenStudy (anonymous):

w:white paint g:green paint w+g=12 16w+20g=12*19=228 solve to get w=3